# Alleviating Forgetfulness of Linear Attention by Hybrid Sparse Attention and Contextualized Learnable Token Eviction

### v1

@@ 3.1.3 Inference implementation
Our models are trained on fixed-length sequences ($N=4096$ in our experiments). Although the number of retained tokens is regularized by the sparsity penalty, the actual computation depends on the retention score of each KV predicted from the input, and no hard cap is enforced as mentioned above. The case is different during inference as we expect guaranteed time and space complexity given arbitrary input with arbitrary length. Therefore we still have to resort to a global budget cap: the number of KVs out of SWA are restricted to at most $b$ per head.
@@
We do not need to compute $r_{j}$ immediately when the computation is possible (i.e. current $n$ reaches $j+R//2$. Instead, we can defer and batchify the LTE computation as long as all the tokens within the receptive field remain in-window. Therefore, we trigger LTE scoring only when uncomputed LTE score $r_{j}$ requires a token that is about to leave the window, i.e. Nw=jR//2N-w=j-R//2. We then rotate the buffer to move next_ptr to 0, and compute the LTE score of tokens from Nw+R//2N-w+R//2 to NR//2N-R//2 altogether. Another problem is that LTE consumes pre-RoPE keys, while KV cache stores post-RoPE keys, hence we need to undo RoPE on cached keys by applying the corresponding inverse rotation. With RoPE implemented using cached sinusoids, this is equivalent to applying RoPE on the keys with negated sine values at offset NwN-w. With this trick, LTE scoring is carried out only every WRW-R steps in parallel with minimal overhead.
@@ 4 Empirical studies
We evaluate hybrid linear-attention models that interleave Gated DeltaNet (GDN) with various mixers from the hierarchy in Figure 1: sliding-window attention (GDN+SWA), learnable token eviction (laLTE), native sparse attention (laNSA), full-attention (GDN+Attn.), plus several ablated variants of LTE. We compare against pure GDN, NSA, and a strong transformer (Transf.++) following Gu & Dao (2024). Given limited computational resources available, we train 0.4B- and 1.4B-sized models on 10B and 30B tokens from FineWeb-Edu (Penedo et al., 2024), following Yang et al. (2024b). We implement LTE layers, hybrid models, as well as inference kernels for LTE and NSA on top of Flash Linear Attention (Yang & Zhang, 2024). Below we leverage both short-context and long-context/retrieval-intensive benchmarks to assess models above and to demonstrate the effectiveness of our laLTE and laNSA approaches, and compare the prefilling and decoding efficiency. More details are available in Appendix A, and we further inspect the retention rates in Appendix B.
@@ 4.1 Short-context benchmarks
Table 1: Results (perplexity and accuracy%) on short-context tasks. Highest numbers are bolded.
@@ 1.4B Params, 30B Tokens
We first evaluate zero-shot commonsense tasks using LM-evaluation-harness (Gao et al., 2024). Given short inputs (generally shorter than the SWA window), we would not expect full attention to confer an advantage. Thus performance should be broadly similar across models, and differences most likely reflect training quality rather than access to distant tokens. This is confirmed by results in Table 1. At 1.4B, GDN slightly outperforms Transformer++, while NSA and GDN+Attn. fall in between; the ordering is not consistent at 0.4B. GDN+SWA performs similar or slightly better, while adding LTE or NSA further yields small gains. Overall, the gaps are small and can only corroborate Yang et al. (2024b)’s observations that linear attention and hybrid models are comparable to or slightly better than transformers on short-context tasks, including our laLTE and laNSA.
@@
Results in Table 2 and Table 3 confirm the effectiveness of hybrid models, laLTE and laNSA in particular. Specifically, pure GDN excels on S-NIAH; at 1.4B it approaches full transformers, echoing Yang et al. (2024b). However, GDN lags on EVAPORATE. GDN+SWA considerably improves on it, but is much weaker on S-NIAH, especially beyond its 1K window size and at 1.4B. At the cost of efficiency, GDN+Attn. consistently achieves best EVAPORATE results across scales, surpassing full transformers, and performs comparably to full transformers and pure GDN on S-NIAH. As for our models, despite similar time and space complexity, laLTE outperforms pure GDN and GDN+SWA on both suites, and approaches GDN+Attn. on S-NIAH at 1.4B. This confirms laLTE’s effectiveness and versatility for long-context retrieval under strict complexity budgets. With full history kept in memory, laNSA further attains the best results among linear-time models on EVAPORATE, approaches full transformers on 0.4B scale S-NIAH, and outperforms pure NSA. However it trails full transformer and GDN+Attn. on EVAPORATE and weaker on 1.4B S-NIAH, especially with longer contexts despite theoretical all-history access. Notably, NSA requires at least 16 group size for grouped-query attention. In our models with relatively small hidden sizes and head counts, NSA is rendered closer to multi-query attention, which can be a performance bottleneck.
@@
We further ablate laLTE against GDN+following token mixers with constant time and space complexity: LTE-MLP replaces CNN with a 2-layer MLP with no context access (especially no future access); Pure LTE is a pure transformer with all layers sparsified by LTE; Unif.+SWA retains out-of-window tokens at a uniform stride; TOVA is one of the state-of-the-art eviction heuristics based on accumulated attention scores (Oren et al., 2024); the latter two are training-free. Although LTE-MLP, and TOVA achieve moderate results on EVAPORATE, better than pure GDN, all of them underperform laLTE, especially on S-NIAH; while Uniform performs much worse than all other models. This confirms the importance of an accurate eviction rule, and underscores the effectiveness of our design of contextualized and learnable eviction. The observation is similar on Pure LTE, which supports our choice to combine sparse attention with linear attention to compensate for the evicted memory. To summarize, all our results confirm the importance of direct and accurate access of past tokens on retrieval-intensive tasks. From GDN, GDN+SWA, laLTE, laNSA, to GDN+Attn., more complete direct access to past tokens generally lead to better performance, though task-level outcomes vary; while laLTE proves its superiority for accurate access. With constrained time and space complexity, the results highlight the effectiveness of our laLTE and laNSA methods.
@@ 4.3 Efficiency
Figure 4: Comparing time costs per token between LTE and SWA or full attention with Flash Attention-2, on (a) kernels with an artificial KV cache during prefilling; and hybrid layers with real inputs during (b) prefilling and (c) decoding, using our 1.4B models with batch size=32.
We then benchmark the actual computational efficiency of our LTE approach. All results are measured on a single H100 GPU following our 1.4B configuration with batch size=32. Figure 4 (a) shows the runtime of prefilling kernels at different sequence lengths, compared with SWA and full attention using the highly-optimized FlashAttention-2 (Dao, 2024) kernels. As for LTE we use an artificial half-full out-of-window KV-cache, with 1024 tokens identical to the SWA window size. Our SWA+KV-sparse attention kernel exhibits a linear-growing runtime comparable to SWA, and much faster than full attention. While LTE depends on real context for token eviction decision; many layers and heads retain only a few tokens, but the LTE scoring and caching introduce some additional overheads. We therefore also measure the prefilling and decoding time of our actual 1.4B models on real inputs, implemented based on Flash Linear Attention. To better isolate the time costs of sequence mixers, we only measure the time of hybrid layers. The results similarly confirm the efficiency of our LTE implementation, as shown by Figure 4 (b)-(c). It is further noteworthy that LTE uses a substantially smaller and constant-size KV-cache compared to full attention, enabling much larger batch sizes and token throughput. Taken together, these results support adopting LTE in lieu of SWA for better performance without degrading efficiency.
@@ Appendix A Implementation details
We train our models using Flame (Zhang & Yang, 2025) and refer to their training recipes, while adapted to our computational resources available. All the models are implemented based on Flash Linear Attention (FLA). Particularly, FLA implements the prefilling kernel for NSA, and we develop the decoding kernel based on it. Both the 0.4B and 1.4B models have 24 layers, with hidden size = 1024 and 2048 respectively. As for GDN layers, we use 8 heads and expand_v = 1. As for attention layers, we set number of heads $H_{q}=H_{kv}=16$ at 0.4B, and grouped-query attention (Ainslie et al., 2023a) with $H_{q}=32,H_{kv}=8$ at 1.4B. Similar configurations are also adopted by Wang et al. (2025). The exception is that NSA requires at least 16 group size due to hardware constraints, hence we use $H_{q}=32,H_{kv}=2$. Token embeddings are tied on 0.4B models. The LTE module consists of a 3-layer 1D grouped convolution stack, using kernel size = 3 and dilation = 2, yielding a receptive field $R=13$. Each layer halves the channel width. Each layer is followed by a Swish activation and a dropout layer with $p=0.1$. It then passes linear projection per head implemented as grouped convolution with kernel size = 1, and produces the retention score $r$ after sigmoid. We implement directly with Conv2D to avoid tensor manipulation overheads. The number of groups equals the number of KV heads to process each head independently and efficiently. The LTE module will add $\sim$1.7% and $\sim$0.7% parameters on 0.4B and 1.4B models respectively.
We then train models by AdamW optimizer with learning rate = 3e-4 and a cosine decay schedule. At 0.4B/1.4B, we use global batch sizes of 0.5M/1.0M tokens, context length 4096, for 20480/30720 steps with 1024/512 warmup steps. As for LTE training, we initialize the regularization weight matrix $\lambda$ to 1e-9, and update it via a feedback loop every $u=32$ steps for stability, following the pseudocode given in Algorithm 1. The update is based on the exponential moving average $\bar{c}$ of the retention count $c$, with coefficient $\alpha_{ema}=2/(1+u/2)$ to approximate the average within the update cycle.
@@
As for evaluation, we use LM-eval-harness for short-context and RULER evaluations, and leverage the reference implementation by Arora et al. (2024c) for EVAPORATE benchmarks. We follow the EVAPORATE input truncation scheme, but only truncate the input length to 4K tokens when it exceeds 4K to better reflect long-context performance. Inference is carried out using greedy decoding in FP32 for stability, except that GDN and Flash Attention kernels support FP16 only, hence we perform all token-mixing ops in FP16 for parity. Training-free sparse attention mechanisms including Unif.+SWA and TOVA are evaluated by applying them to pretrained GDN + Attn. We use an identical 1024 total cache size. For compute-cost measurements, we report the total elapsed time between the CUDA event pairs at layer entry after a full warm-up phase; for prefill we average the middle three of five runs. We use samples from the training dataset as inputs.
@@ Appendix B Retention pattern
Figure 5 shows the average retention rate per head per layer from the out-of-SWA part of 16 samples drawn from the training set produced by our l.4B laLTE model. We find that retention rates are highly varied across layers and heads, while the higher layers are denser. This is similar to findings in other works, e.g. Łańcucki et al. (2025), except that dense heads do not present in lowest layers; The interleaved GDN layers may have a role in it. The observation also supports the necessity to have fine-grained per-head cache budget.

### v2

@@ 3.1.3 Inference implementation
Our models are trained on fixed-length sequences ($N=4096$ in our experiments). Although the number of retained tokens is regularized by the sparsity penalty, the actual computation depends on the retention score of each KV predicted from the input, and no hard cap is enforced as mentioned above. The case is different during inference as we expect guaranteed time and space complexity given arbitrary input with arbitrary length. Therefore we still have to resort to a global budget cap: the number of cached KVs out of window is restricted to at most $b$ per head.
@@
We do not need to compute $r_{j}$ immediately when the computation is possible (i.e. current $n$ reaches $j+R//2$. Instead, we can defer and batchify the LTE computation as long as all the tokens within the receptive field remain in-window. Therefore, we trigger LTE scoring only when uncomputed LTE score $r_{j}$ requires a token that is about to leave the window, i.e. Nw=jR//2N-w=j-R//2. We then rotate the buffer to move next_ptr to 0, and compute the LTE score of tokens from Nw+R//2N-w+R//2 to NR//2N-R//2 altogether. Another problem is that LTE consumes pre-RoPE keys, while KV cache stores post-RoPE keys, hence we need to undo RoPE on cached keys by applying the corresponding inverse rotation. With RoPE implemented using cached sinusoids, this is equivalent to applying RoPE on the keys with negated sine values at offset NwN-w. With this trick, LTE scoring is carried out only every WRW-R steps in parallel with minimal amortized overhead.
@@ 4 Empirical studies
We evaluate hybrid linear-attention models that interleave Gated DeltaNet (GDN) with various mixers from the hierarchy in Figure 1: sliding-window attention (GDN+SWA), learnable token eviction (laLTE), native sparse attention (laNSA), full-attention (GDN+Attn.), plus several ablated variants of LTE. We compare against pure GDN, NSA, and a strong transformer (Transf.++) following Gu & Dao (2024). Given limited computational resources available, we train 0.4B- and 1.4B-sized models on 10B and 30B tokens from FineWeb-Edu (Penedo et al., 2024), following Yang et al. (2024b). We implement LTE layers, hybrid models, as well as inference kernels for LTE and NSA on top of Flash Linear Attention (Yang & Zhang, 2024). Below we leverage both short-context and long-context/retrieval-intensive benchmarks to assess models above and to demonstrate the effectiveness of our laLTE and laNSA approaches, and compare the prefilling and decoding efficiency. More details are available in Appendix A. Additional results on KV retention rates are given in Appendix B.
@@ 4.1 Short-context benchmarks
Table 1: Results (perplexity and accuracy%) on short-context tasks. Best numbers are bolded.
@@ 1.4B Params, 30B Tokens
We first evaluate zero-shot commonsense tasks using LM-evaluation-harness (Gao et al., 2024). Given short inputs (generally shorter than the SWA window), we do not expect denser attention to confer any advantage. Thus performance should be broadly similar across models, and differences most likely reflect training quality rather than access to distant tokens. This is confirmed by results in Table 1. At 1.4B, GDN slightly outperforms Transformer++, while NSA and GDN+Attn. fall in between; the ordering is not consistent at 0.4B. GDN+SWA performs similar or slightly better, while adding LTE or NSA further yields small gains. Overall, the gaps are small and can only corroborate Yang et al. (2024b)’s observations that linear attention and hybrid models (including our laLTE and laNSA) are comparable to or slightly better than transformers on short-context tasks.
@@
Results in Table 2 and Table 3 confirm the effectiveness of hybrid models, laLTE and laNSA in particular. Specifically, pure GDN excels on S-NIAH; at 1.4B it approaches full transformers, echoing Yang et al. (2024b). However, GDN lags on EVAPORATE. GDN+SWA considerably improves on it, but is much weaker on S-NIAH, especially beyond its 1K window size and at 1.4B. At the cost of efficiency, GDN+Attn. consistently achieves best EVAPORATE results across scales, surpassing full transformers, and performs comparably to full transformers and pure GDN on S-NIAH. As for our models, despite similar time and space complexity, laLTE outperforms pure GDN and GDN+SWA on both suites, and approaches GDN+Attn. on S-NIAH at 1.4B. This confirms laLTE’s effectiveness and versatility for long-context retrieval under strict complexity budgets. With full history kept in memory, laNSA further attains the best results among linear-time models on EVAPORATE, approaches full transformers on 0.4B scale S-NIAH, and outperforms pure NSA. However it trails full transformer and GDN+Attn. on EVAPORATE and weaker on 1.4B S-NIAH, despite theoretical all-history access. Notably, NSA requires at least 16 group size for grouped-query attention. In our models with relatively small hidden sizes and head counts, NSA is rendered closer to multi-query attention, which can be a performance bottleneck and a factor of instability.
@@
We further ablate laLTE against GDN+following token mixers with constant time and space complexity: LTE-MLP replaces CNN with a 2-layer MLP with no context access (especially no future access); Pure LTE is a pure transformer with all layers sparsified by LTE; Unif.+SWA retains out-of-window tokens at a uniform stride; TOVA is one of the state-of-the-art eviction heuristics based on accumulated attention scores (Oren et al., 2024); the latter two are training-free. Although LTE-MLP and TOVA achieve moderate results on EVAPORATE, better than pure GDN, all of them underperform laLTE, especially on S-NIAH; while Unif.+SWA performs much worse than all other models. This confirms the importance of an accurate eviction rule, and underscores the effectiveness of our design of contextualized and learnable eviction. The observation is similar on Pure LTE, which supports our choice to combine sparse attention with linear attention to compensate for the evicted memory. To summarize, all our results confirm the importance of direct and accurate access of past tokens on retrieval-intensive tasks. From GDN, GDN+SWA/laLTE, laNSA, to GDN+Attn., more complete direct access to past tokens generally allows better performance, though task-level outcomes vary; while our laLTE proves its strength for accurate access under rigorous time and space complexity constraints.
@@ 4.3 Efficiency
Figure 4: Comparing time costs between LTE and SWA or full attention with Flash Attention-2, on (a) kernels with an artificial KV cache during prefilling; and hybrid layers with real inputs during (b) prefilling and (c) decoding with a 4K-token prompt, using our 1.4B models with batch size=32.
We then benchmark the actual computational efficiency of our LTE approach. All results are measured on a single H100-SXM GPU following our 1.4B configuration with batch size=32. Figure 4 (a) shows the runtime of prefilling kernels at different sequence lengths, compared with SWA and full attention using the highly-optimized FlashAttention-2 (Dao, 2024) kernels. As for LTE we use an artificial half-full out-of-window KV-cache, with 1024 tokens identical to the SWA window size. Our SWA+KV-sparse attention kernel exhibits a linear-growing runtime comparable to SWA, and much faster than full attention. While LTE depends on real context for token eviction decision; many layers and heads retain only a few tokens, but the LTE scoring and caching introduce some additional overheads. We therefore also measure the prefilling and decoding time of our actual 1.4B models on real inputs, implemented based on Flash Linear Attention. To better isolate the speed of our targets, we only measure the time costs of hybrid layers (i.e. excluding other layers like GDN). The results similarly confirm the efficiency of our LTE implementation, as shown by Figure 4 (b)-(c). It is further noteworthy that LTE uses a substantially smaller and constant-size KV-cache compared to full attention, allowing much larger batch sizes and token throughput. Taken together, these results support adopting LTE in lieu of SWA for better performance without degrading efficiency.
@@ Appendix A Implementation details
We train our models using Flame (Zhang & Yang, 2025) and refer to their training recipes, while adapted to our computational resources available. All the models are implemented based on Flash Linear Attention (FLA). Particularly, FLA implements the prefilling kernel for NSA, and we develop the decoding kernel based on it. We also develop the LTE prefilling kernel upon the FLA attention kernel. Both the 0.4B and 1.4B models have 24 layers, with hidden size = 1024 and 2048 respectively. As for GDN layers, we use 8 heads and expand_v = 1. As for attention layers, we set number of heads $H_{q}=H_{kv}=16$ at 0.4B, and grouped-query attention (Ainslie et al., 2023a) with $H_{q}=32,H_{kv}=8$ at 1.4B. Similar configurations are also adopted by Wang et al. (2025). The exception is that NSA requires at least 16 group size due to hardware constraints, hence we use $H_{q}=32,H_{kv}=2$. Token embeddings are tied on 0.4B models. The LTE module consists of a 3-layer 1D grouped convolution stack, using kernel size = 3 and dilation = 2, yielding a receptive field $R=13$. Each layer halves the channel width. Each layer is followed by a Swish activation and a dropout layer with $p=0.1$. It then passes linear projection per head implemented as grouped convolution with kernel size = 1, and produces the retention score $r$ after sigmoid. We implement directly with Conv2D to avoid tensor manipulation overheads. The number of groups equals the number of KV heads to process each head independently and efficiently. The LTE module will add $\sim$1.7% and $\sim$0.7% parameters on 0.4B and 1.4B models respectively.
We use the AdamW optimizer with learning rate = 3e-4 under cosine decay. At 0.4B/1.4B, we use global batch sizes of 0.5M/1.0M tokens, context length 4096, for 20480/30720 steps with 1024/512 warmup steps. As for LTE training, we initialize the regularization weight matrix $\lambda$ to 1e-9, and update it via a feedback loop every $u=32$ steps for stability, following the pseudocode given in Algorithm 1. The update is based on the exponential moving average $\bar{c}$ of the retention count $c$, with coefficient $\alpha_{ema}=2/(1+u/2)$ to approximate the average within the update cycle.
@@
During LTE decoding, we directly use the flash_attention_kv_cache API by Flash Attention, which supports different past KV cache length per sample in a batch by passing in cache_seqlens. Since we have different cache length per head, we merge the dimensions for KV head and batch size, leaving a KV cache with the number of query heads equals to the GQA group size. As for evaluation, we use LM-eval-harness for short-context and RULER evaluations, and leverage the reference implementation by Arora et al. (2024c) for EVAPORATE benchmarks. We follow their input truncation scheme, but only truncate the input length to 4K tokens when it exceeds 4K to better reflect long-context performance. Inference is carried out using greedy decoding in FP32 for stability, except that GDN and Flash Attention kernels support FP16 only, hence we perform all token-mixing ops in FP16 for parity. Training-free sparse attention mechanisms including Unif.+SWA and TOVA are evaluated by applying them to pretrained GDN + Attn. We use an identical 1024 total cache size for both of them. For compute-cost measurements, we report the total elapsed time between the CUDA event pairs at layer entry after a full warm-up phase; for prefilling we average the middle three of five runs. Samples from the training dataset are used as inputs.
@@ Appendix B Retention pattern
Figure 5 shows the average retention rate per head per layer from the out-of-SWA part of 16 samples drawn from the training data produced by our l.4B laLTE model. We find that retention rates are highly varied across layers and heads, while the higher layers are often denser. This is similar to findings in other works, e.g. Łańcucki et al. (2025), except that dense heads do not present in lowest layers; The interleaved GDN layers may have a role in it. The observation also supports the necessity of having a fine-grained cache budget different from head to head and from layer to layer.